OpenStudy (anonymous):

e^x=-4 Solve the equation for x. If necessary, round your answer to two decimal places. Do not include "x =" in your answer.

jimthompson5910 (jim_thompson5910):

e^x = -4 ln(e^x) = ln(-4) Since you cannot take the natural log of a negative number, this means that we can stop and simply say that there are no solutions.

OpenStudy (anonymous):

I wrote the problem wrong it does have an In< Its In e^x=-4

jimthompson5910 (jim_thompson5910):

ln(e^x) = -4 x*ln(e) = -4 x*1 = -4 x = -4 So the solution is x = -4

jhonyy9 (jhonyy9):

the right part not is correct because (-4) using ln can being writing -4lne that we know that lne =1 so than lne^x =xlne - and -4lne =xlne so than x=-4

OpenStudy (anonymous):

e^x is always >0 so it cannot be equal to -4, No Solution.

jhonyy9 (jhonyy9):

how do you think e^(-4) = 1/(e^4) not is always greater than zero ?

OpenStudy (anonymous):

@jhonyy9 you need to review exponential functions. Have you seen the graph of the function e^x. Look at the graph attached and http://archives.math.utk.edu/visual.calculus/0/exp_log.5/index.html

jhonyy9 (jhonyy9):

thank you - yes you are right good luck bye